OpenStudy (anonymous):

A faulty odometer (mileage indicator) on a car registers 6 miles for every 5 miles actually traveled. At the start of a journey it reads 500 miles and the end of the journey it reads 560 miles. What distance has actually been covered?

OpenStudy (anonymous):

6/5 =60/x x=50 50 miles